# Efficient Dual-domain Image Dehazing with Haze Prior Perception

### v1

@@ III-B1 Motivation
Global modeling ensures structural consistency, while local modeling restores fine textures essential for high-quality recovery. Some Transformer-based methods [41, 42] enhance local perception using single-scale depth-wise convolutions, but often neglect multi-scale feature learning, which is crucial for handling uneven haze.
To tackle this, we propose MGAM, which enhances local modeling through multi-scale feature extraction. By stacking small-kernel convolutions, MGAM captures multi-level representations and expands the receptive field at low computational cost. Unlike DSANet [43], which uses parallel large-kernel convolutions, our approach extracts features across different scales more efficiently. Here, “level” borrows the convention from U-Net, referring to receptive field size rather than spatial resolution.
Moreover, we introduce a hybrid gating mechanism that leverages low-level features rich in edge and texture cues to regulate the propagation of high-level semantic information. This modulation enables hierarchical interaction, facilitating adaptive feature fusion while preserving fine details. Additionally, the sensitivity of low-level features to local haze variations enhances the model’s ability to maintain semantic consistency and robustness under non-uniform haze conditions.
@@ III-B2 Pipeline
Given the output feature $\hat{X}_{f}^{i,j}$ of HAFM and the intermediate dark channel-guided feature $\hat{X}_{d}^{i,j}$, the processing flow of MGAM is defined as:
@@
where $X_{f}^{i,j}$ denotes the residual output feature of MGAM, and $\tilde{X}_{d}^{i,j}$ is the dark channel correction feature.
In MGAM, we first apply a $1\times 1$ convolution to the input feature (after BN) to double the number of channels. As shown in Fig. 3 (c), the resulting feature is split into two parts: one for extracting multi-scale features $X_{fea}^{i,j}$, and the other for generating gating signals $X_{gate}^{i,j}$. This process is expressed as:
@@
Multi-scale Feature Extraction. Each of the two parallel branches applies two stacked depth-wise convolutions for multi-scale feature extraction. The first convolution in each branch uses $k\times k$ kernels, while the second uses the same kernels with a dilation factor of 2 to expand the receptive field. The process is defined as follows:
$X_{fea_{k}}^{i,j}=\mathrm{DDConv}_{k\times k}(\mathrm{DConv}_{k\times k}(X_{fea}^{i,j})),\quad k\in\{3,5\},$ (11)
where $X_{fea_{k}}^{i,j}$ represents the output feature of the $k\times k$ branch.
Gating Signals Generation. A dual-branch structure generates gating signals, where each branch applies a depth-wise convolution with $k\times k$ kernels and a Sigmoid activation:
$X_{gate_{k}}^{i,j}=\mathrm{Sigmoid}(\mathrm{DConv}_{k\times k}(X_{gate}^{i,j})),\quad k\in\{3,5\},$ (12)
where $X_{gate_{k}}^{i,j}$ represents the gating signal corresponding to the $k\times k$ feature branch.
Multi-scale Feature Fusion. The gated features are obtained via element-wise multiplication of $X_{fea_{k}}^{i,j}$ and $X_{gate_{k}}^{i,j}$:
$X_{gated_{k}}^{i,j}=X_{fea_{k}}^{i,j}\odot X_{gate_{k}}^{i,j},\quad k\in\{3,5\}.$ (13)
These features are concatenated and processed by a $1\times 1$ convolution with GELU activation, producing the fusion feature $X_{dual}^{i,j}$. To mitigate early-stage gating instability, a skip connection incorporates the original input feature before a final $1\times 1$ convolution and GELU activation, yielding the multi-scale fusion feature $X_{mult}^{i,j}$. The entire process is formulated as follows:
@@
Feature Feedback Module. A Channel Attention Module [44] (CAM) is applied at the end of MGAM to generate feedback for PCGB while reducing redundancy in multi-scale feature extraction. The attention map $M_{ca}$ is computed from $X_{mult}^{i,j}$ and applied element-wise to both $X_{mult}^{i,j}$ and $\hat{X}_{d}^{i,j}$:
@@
$\tilde{X}_{d}^{i,j}=\mathrm{Conv}_{1\times 1}(\hat{X}_{d}^{i,j}\odot M_{ca}),$ (17)
where $X_{f}^{i,j}$ represents the final output and $\tilde{X}_{d}^{i,j}$ is the feedback dark channel correction feature. This operation enables PCGB to utilize the information of $X_{mult}^{i,j}$ while selectively emphasizing important channels, enhancing its connection with the dehazing branch.
@@ III-C1 Motivation
Although the dark channel prior provides a rough estimation of haze distribution, it often fails in outdoor scenes containing sky regions. Its assumption of locally uniform haze does not hold in real-world scenarios with complex backgrounds and varying haze densities, limiting its effectiveness for accurate haze localization.
Still, physically grounded priors provide useful inductive bias, which helps models generalize better to unseen conditions. For instance, Dehamer [35] incorporates the dark channel prior into positional encodings to enhance haze perception. However, it treats the prior as static and lacks correction during inference, which may lead to suboptimal guidance when the prior is inaccurate.
To address this, we propose the Prior Correction Guidance Branch (PCGB). Instead of using the dark channel prior as a fixed input, PCGB treats it as a haze-awareness cue and progressively refines the guidance features through feedback from the main dehazing branch. This design preserves the benefits of inductive bias while improving haze localization under complex and non-uniform conditions.
@@ III-C2 Pipeline
As illustrated in Fig. 4, given the dark channel $X_{d}\in\mathbb{R}^{H\times W}$ of the hazy image $I$, we first encode the features using two successive convolutional layers:
@@
where $X_{d}^{1}\in\mathbb{R}^{C\times H\times W}$ denotes the initial dark channel feature of the first stage. Then, downsampling is performed to derive the initial dark channel features for the second and third stages:
$X_{d}^{i}=f_{dw}(X_{d}^{i-1}),\quad i\in\{2,3\}$ (19)
where $X_{d}^{i}$ denotes the initial dark channel feature for the $i$-th stage, and $f_{dw}$ denotes downsampling, which halves the spatial resolution while doubling the channel number.
Dark Channel-Guided Features in the Encoder. The first DGFDBlock in the first stage uses $X_{d}^{1}$ as the dark channel-guided feature, denoted as $X_{d}^{1,0}=X_{d}^{1}$. In later stages, the first DGFDBlock obtains its dark channel-guided feature by downsampling the feedback dark channel correction from the last DGFDBlock of the previous stage and fusing it with the initial dark channel feature of the current stage via SKFusion [14]. For subsequent DGFDBlocks within a stage, the feedback dark channel correction from the previous DGFDBlock is directly fused with the initial dark channel feature. The process is expressed as:
@@
Here, $X_{d}^{i,j-1}$ represents the dark channel-guided feature for the $j$-th DGFDBlock at the $i$-th stage in the encoder. $N_{i-1}$ is the number of DGFDBlocks in the previous stage, and $\tilde{X}_{d}^{i,j-1}$ denotes the feedback dark channel correction from the preceding DGFDBlock. $\mathrm{SKF}_{i}$ is the SKFusion [14] block shared across all DGFDBlocks within the $i$-th stage. The downsampling operation $f_{dw}$ maintains parameter consistency with the one used for initial dark channel features.
Dark Channel-Guided Features in the Decoder. The decoder follows a reversed process, propagating from Stage 3 back to Stage 1. Unlike the encoder, the first DGFDBlock in each stage obtains its dark channel-guided feature by upsampling the feedback dark channel correction from the last DGFDBlock of the previous stage and fusing it with the initial dark channel feature of the current stage via SKFusion [14]:
$X_{d}^{i,j-1}=\begin{cases}\mathrm{SKF}_{i}(f_{up}(\tilde{X}_{d}^{i+1,N_{i+1}}),X_{d}^{i}),\quad j=1&\\ \mathrm{SKF}_{i}(\tilde{X}_{d}^{i,j-1},X_{d}^{i}).\quad\quad\quad 1<j\leq N_{i}&\end{cases}$ (21)
Here, $X_{d}^{i,j-1}$ is the dark channel-guided feature for the $j$-th DGFDBlock at the $i$-th stage in the decoder. $N_{i+1}$ denotes the number of DGFDBlocks in the next stage, and $f_{up}$ is an upsampling operation combining a $1\times 1$ convolution with pixel-shuffle, doubling the spatial dimensions and halving the channel number.
III-D Loss Function
To simultaneously improve detail recovery and global structure restoration, we adopt a dual-domain $L_{1}$ loss that supervises both spatial and frequency domains. This encourages the network to balance fine-grained texture reconstruction with consistent global appearance. The overall loss is defined as:
where $I_{c}$ is the dehazed output of DGFDNet, $I_{g}$ is the ground-truth image, $\mathcal{F}$ denotes the Fast Fourier Transform, and $\lambda$ is a balancing parameter that controls the contribution of the frequency domain loss. In our experiments, $\lambda$ is set to $0.1$.
IV Experiments
IV-A Experimental settings
Datasets. We evaluate the proposed DGFDNet on both synthetic and real-world datasets. The synthetic dataset is the RESIDE [45] dataset, which includes two training sets: the ITS with 13,990 indoor image pairs and the OTS with 313,950 outdoor image pairs. The test set SOTS consists of 500 indoor and 500 outdoor image pairs. For real-world datasets, we use Dense-Haze [46] and NH-HAZE [47], each containing 55 paired real images. The last five pairs of each dataset are used for testing, and the rest for training.
TABLE I: Detailed training configuration of different evaluation datasets.
Evaluation dataset Init lr Batch size Patch size Epochs
SOTS-Indoor 3e-4 8 $256\times 256$ 1000
SOTS-Outdoor 1e-4 8 $256\times 256$ 60
Dense-Haze 2e-4 2 $512\times 512$ 5000
NH-HAZE 3e-4 4 $384\times 384$ 6000
Comparison settings. We compare DGFDNet with ten CNN-based methods: GridDehazeNet [24], MSBDN [48], FFA-Net [25], PMNet [49], MAXIM-2S [50], DEA-Net [28], FocalNet [40], OKNet [51], PGH2Net [52], and DCMPNet [53]. We further include three Transformer-based models: DeHamer [35], DehazeFormer [14], and Fourmer [54]. All methods are evaluated using Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index Measure (SSIM).
TABLE II: Quantitative comparisons with state-of-the-art dehazing methods on the synthetic and real-world datasets.
@@ IV-A Experimental settings
Training details. Each stage of the model contains 2, 2, 4, 2, 2 DGFDBlocks, with a base channel size of 32. We use the Adam optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The initial learning rate is dataset-dependent and is gradually reduced to $1\times 10^{-6}$ following a cosine annealing schedule (Table I). Data augmentation includes random cropping and flipping. FLOPs are calculated on $256\times 256$ image patches. All experiments are conducted on an NVIDIA 3090 GPU.
@@ IV-B1 Quantitative Comparisons
Table II presents quantitative comparisons between DGFDNet and state-of-the-art methods on two synthetic and two real-world datasets. Bold and underlined values indicate the best and second-best results, respectively. Overall, our method outperforms all existing approaches across all datasets.
Compared to classical CNN-based methods like FocalNet [40] and OKNet [51], DGFDNet improves PSNR by 1.36 dB and 1.39 dB on SOTS-Indoor, and by 0.80 dB and 0.83 dB on SOTS-Outdoor, while using roughly half the parameters and less than half the FLOPs. Unlike FocalNet, which also adopts dual-domain selection mechanism with spatial attention, our method integrates dark channel-guided attention for more precise haze localization and a full-frequency modulation unit to better handle varying degradation levels.
Compared to the heavy DCMPNet [53], DGFDNet achieves comparable results on SOTS-Indoor with only 12.0% of its parameters and 19.7% of its FLOPs. On SOTS-Outdoor, DGFDNet surpasses DCMPNet by 1.95 dB PSNR and 0.002 SSIM, demonstrating enhanced robustness in complex outdoor conditions. This improvement mainly arises from the PCGB module, which effectively mitigates the limitations of traditional dark channel priors.
Under similar resource budgets, DGFDNet consistently outperforms the lightweight PGH2Net [52], achieving 0.48 dB higher PSNR on SOTS-Indoor and 0.99 dB PSNR plus 0.006 SSIM gains on SOTS-Outdoor, indicating a superior trade-off between accuracy and efficiency.
Beyond synthetic datasets, our method performs exceptionally well on real-world datasets, including Dense-Haze [46] and NH-HAZE [47]. While its PSNR on NH-HAZE is slightly lower than Dehamer [35], it ranks second overall. Notably, it achieves the highest SSIM on both datasets, demonstrating its strong generalization and ability to reconstruct global structures in real-world hazy scenarios.
@@ IV-B2 Qualitative Comparisons
Fig.5 shows dehazing results on synthetic images from the SOTS dataset. Our method consistently achieves the highest PSNR across all testing images. In the first row of Fig. 5(b, c), FFA-Net [25] and MAXIM-2S [50] produce blurred edges and noticeable artifacts in complex indoor scene. Additionally, MAXIM-2S [50] and Dehamer [35] exhibit insufficient haze removal in deeper regions, as shown in the second row of Fig. 5(c, d). While FocalNet [40] and OKNet [51] perform better, our method still surpasses them. As shown in Fig. 5(f), our approach not only removes haze more effectively but also preserves sharp edges, delivering superior visual quality. In outdoor scenes, our method continues to outperform others, achieving more thorough haze removal and clearer reconstruction.
Fig.6 and Fig.7 show real-world dehazing results on Dense-Haze [46] and NH-HAZE [47] datasets. In Fig.6, two dense haze scenes are compared. In the first scene, our method achieves slightly lower PSNR than Dehamer[35] but surpasses it by 0.183 in SSIM, delivering clearer textures and reduced color distortion. The second scene features severe haze with significant detail loss. Even in this challenging case, our method better restores the overall structure of the image and preserves more details compared to Dehamer [35].
Fig. 7 compares two non-homogeneous haze scenes. Except for a slightly lower PSNR in the second scene compared to FocalNet[40], our method leads on all other metrics. In the first scene, our dehazing result exhibits minimal color distortion and avoids the heavy artifacts introduced by other methods. In the second scene, our method removes haze most thoroughly, revealing the clearest details, though it also mistakenly removes some snow traces on the ground.
@@
IV-C Ablation Studies
We first evaluate the individual and combined contributions of HAFM, MGAM, and PCGB to verify their effectiveness and complementarity. Next, we perform ablation studies by modifying each module independently while keeping the rest of the architecture fixed. All models are trained on the ITS dataset and evaluated on SOTS-Indoor under the same conditions as the final model. Since PCGB targets the limitations of dark channel priors in outdoor scenes, we additionally validate its performance on SOTS-Outdoor using a model trained on the OTS dataset. Finally, attention map visualizations are used to demonstrate the corrective effect of PCGB.
TABLE III: Ablation study for three core modules.
@@ IV-C Ablation Studies
9.79 dB 32.69 dB 34.27 dB 32.18 dB 33.67 dB 32.58 dB 35.20 dB PSNR
11.54 dB 34.49 dB 30.01 dB 34.38 dB 40.42 dB 40.29 dB 41.06 dB PSNR
11.10 dB 31.53 dB 32.51 dB 33.70 dB 35.50 dB 37.10 dB 38.36 dB PSNR
22.04 dB 31.72 dB 29.04 dB 28.88 dB 35.11 dB 35.82 dB 39.72 dB PSNR
12.35 dB 26.62 dB 28.41 dB 26.86 dB 30.09 dB 31.83 dB 33.39 dB PSNR
Figure 5: Visual comparisons on synthetic hazy images from the SOTS dataset. Key regions highlighted by red boxes are enlarged and placed in the lower-left corner of each image for clearer comparison.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
(a) Hazy image (b) Dehamer [35] (c) Ours (d) GT
Figure 6: Visual comparisons on real-world hazy images from the Dense-Haze datasets. Key regions marked with red boxes are enlarged in left-to-right order and arranged vertically on the right.
(a) Hazy image (b) Dehamer [35] (c) FocalNet [40] (d) Ours (e) GT
IV-C1 Effectiveness of Core Modules
To thoroughly evaluate the effectiveness and complementarity of the three core modules, we design four variants: MGAM, HAFM (without feedback correction), HAFM+PCGB (with complete feedback-guided correction), and MGAM+HAFM. Table III summarizes the performance of each variant.
As shown in Table III, MGAM improves PSNR by 0.75 dB over HAFM with the same SSIM of 0.994, highlighting its advantage in local detail enhancement. Introducing PCGB into HAFM yields an additional 0.70 dB gain, demonstrating its role in enhancing global degradation modeling through adaptive prior correction. Combining MGAM and HAFM brings a substantial PSNR increase of 1.80 dB over MGAM alone and 2.55 dB over HAFM. This gain benefits from the synergy between multi-scale spatial feature extraction and dual-domain modulation. Further adding PCGB achieves the best overall result, confirming that the three modules are complementary and jointly contribute to both local and global restoration. These improvements are achieved with minimal additional computational cost.
IV-C2 Effectiveness of HAFM
To validate HAFM, we design four variants: HAFM-S (spatial modulation only), HAFM-F (frequency modulation only), HAFM-SSA (standard spatial attention replacing dark channel-guided attention), and HAFM-SMLP (spatial MLP replacing frequency modulation). The results are exhibited in Table IV.
TABLE IV: Ablation study for key components in HAFM.

### v2

@@ III-B1 Motivation
While HAFM effectively captures global degradation, it has limited capacity for modeling fine-detail, which is crucial for high-quality restoration. To remedy this, we propose MGAM to refine HAFM-modulated features and enhance local representation.
Recent methods [33, 18] focus on multi-scale structures to handle non-uniform haze by capturing spatial patterns at various scales. MGAM achieves this by stacking small convolutions that efficiently expand the receptive field while maintaining computational efficiency, in contrast to the large-kernel convolutions used in DSANet [39].
In addition, MGAM incorporates a hybrid gating mechanism that uses low-level features, rich in edges and textures, to regulate the flow of high-level semantic information. This hierarchical interaction enables adaptive feature fusion while preserving fine details. The sensitivity of low-level features to local haze variations further improves semantic consistency and robustness under non-uniform haze conditions.
Furthermore, the refined features from MGAM adjust the haze confidence map of the current DGFDBlock, which is then fed back to the PCGB to adaptively update the dark channel-guided features, enhancing haze localization.
@@ III-B2 Pipeline
Given the output feature $\hat{X}_{f}^{i,j}$ from HAFM and the intermediate dark channel-guided feature $\hat{X}_{d}^{i,j}$, MGAM processes them as follows:
@@
where $X_{f}^{i,j}$ is the residual output of MGAM, and $\tilde{X}_{d}^{i,j}$ is the dark channel correction feature.
Multi-scale Feature Generation. As shown in Fig. 3 (c), MGAM consists of two parallel branches with gating mechanisms operating at different scales. The input feature (after BN) is passed through a $1\times 1$ convolution to double the channel count, then split into two parts: one for feature extraction ($X_{fea}^{i,j}$) and the other for gating signals ($X_{gate}^{i,j}$):
@@
Each branch applies two depth-wise convolutions for multi-scale feature extraction. The first convolution uses $k\times k$ kernels ($k\in{3,5}$), followed by a dilated convolution (dilation factor of 2) with the same kernel size to expand the receptive field, producing multi-scale features $X_{fea_{k}}^{i,j}$. Meanwhile, gating signals $X_{gate_{k}}^{i,j}$ are generated via depth-wise convolution with $k\times k$ kernels and a Sigmoid activation. The gated multi-scale features $X_{gated_{k}}^{i,j}$ are obtained by element-wise multiplication:
$X_{fea_{k}}^{i,j}=\mathrm{DDConv}_{k\times k}(\mathrm{DConv}_{k\times k}(X_{fea}^{i,j})),$ (11)
$X_{gate_{k}}^{i,j}=\mathrm{Sigmoid}(\mathrm{DConv}_{k\times k}(X_{gate}^{i,j})),$ (12)
$X_{gated_{k}}^{i,j}=X_{fea_{k}}^{i,j}\odot X_{gate_{k}}^{i,j}.$ (13)
Multi-scale Feature Fusion. The gated features are concatenated and passed through a $1\times 1$ convolution with GELU activation to produce the fused feature $X_{dual}^{i,j}$. To mitigate early-stage gating instability, a skip connection with the original input is added before a final $1\times 1$ convolution and GELU activation, yielding the multi-scale fusion feature $X_{mult}^{i,j}$:
@@
Feature Feedback Module. At the end of MGAM, a CAM [40] is used to generate feedback for PCGB, helping reduce redundancy in multi-scale feature extraction. The attention map $M_{ca}$ is computed from $X_{mult}^{i,j}$ and applied element-wise multiplication to both $X_{mult}^{i,j}$ and $\hat{X}_{d}^{i,j}$:
@@
$\tilde{X}_{d}^{i,j}=\mathrm{Conv}_{1\times 1}(\hat{X}_{d}^{i,j}\odot M_{ca}).$ (17)
Here, $X_{f}^{i,j}$ is the final output, and $\tilde{X}_{d}^{i,j}$ is the feedback correction of dark channel feature. This operation allows PCGB to refine dark channel features based on channel information from $X_{mult}^{i,j}$, enhancing its alignment with the dehazing branch.
@@ III-C1 Motivation
The dark channel prior often fails in outdoor scenes with sky regions, where naturally low intensities lead to inaccurate transmission estimation. Moreover, its assumption of locally uniform haze does not hold in real-world conditions with complex backgrounds and varying haze densities, limiting its accuracy in haze localization. Despite these limitations, physically grounded priors offer useful inductive bias, allowing better generalization to unseen conditions. For instance, Dehamer [34] integrates the dark channel prior into positional encodings to enhance haze perception, but it treats the prior as static and lacks correction during inference, which can lead to suboptimal guidance when the prior is inaccurate.
To resolve this, we propose the Prior Correction Guidance Branch (PCGB). Instead of using the dark channel prior as a fixed input, PCGB treats it as a dynamic haze-awareness cue, refining guidance features through feedback from the dehazing branch. This preserves the inductive bias benefits while enhancing haze localization in complex, uneven conditions. Furthermore, each refinement output is fused with the initial dark channel features using SKFusion [14], combining self-correction and supervision to prevent error accumulation.
@@ III-C2 Pipeline
As illustrated in Fig. 4, given the dark channel $X_{d}\in\mathbb{R}^{H\times W}$ of the hazy image $I$, we first encode the features using two successive convolutions:
@@
where $X_{d}^{1}\in\mathbb{R}^{C\times H\times W}$ is the initial dark channel feature for the first stage. Downsampling is then applied to generate the initial features for the second and third stages:
$X_{d}^{i}=f_{dw}(X_{d}^{i-1}),\quad i\in\{2,3\}.$ (19)
Dark Channel-Guided Features in the Encoder. In the first stage, the first DGFDBlock uses $X_{d}^{1}$ as the dark channel-guided feature, denoted as $X_{d}^{1,0}=X_{d}^{1}$. For subsequent stages, the first DGFDBlock receives its dark channel-guided feature by downsampling the feedback correction from the last DGFDBlock of the previous stage, and fusing it with the initial dark channel feature of the current stage via SKFusion. For the remaining DGFDBlocks in each stage, the dark channel-guided feature is obtained by fusing the feedback correction from the previous DGFDBlock with the initial dark channel feature of the current stage. The process is defined as:
@@
Here, $X_{d}^{i,j-1}$ represents the dark channel-guided feature for the $j$-th DGFDBlock at the $i$-th stage. $N_{i-1}$ is the number of DGFDBlocks in the previous stage, and $\tilde{X}_{d}^{i,j-1}$ denotes the feedback correction of dark channel from the previous DGFDBlock. $\mathrm{SKF}i$ is the SKFusion block shared across all DGFDBlocks within the $i$-th stage.
Dark Channel-Guided Features in the Decoder. The decoder follows the reverse process, propagating from Stage 3 to Stage 1. Unlike the encoder, the first DGFDBlock in each stage derives its dark channel-guided feature by upsampling the feedback correction of dark channel from the last DGFDBlock of the previous stage and fusing it with the initial dark channel feature of the current stage using SKFusion:
[…]
Here, $X_{d}^{i,j-1}$ is the dark channel-guided feature for the $j$-th DGFDBlock at the $i$-th stage in the decoder. $N_{i+1}$ denotes the number of DGFDBlocks in the next stage.
TABLE I: Quantitative comparisons with state-of-the-art dehazing methods on the synthetic and real-world datasets.
@@
III-D Loss Function
To enhance both fine-grained texture restoration and global appearance consistency, we adopt a dual-domain $L_{1}$ loss that supervises both the spatial and frequency domains. The total loss is defined as:
where $I_{c}$ is the dehazed output of DGFDNet, $I_{g}$ is the ground-truth image, and $\lambda$ is set to $0.1$ to control the balance between the two domain losses, following the setting in [38, 45].
IV Experiments
IV-A Experimental settings
Datasets. We evaluate DGFDNet on synthetic and real-world datasets. The synthetic RESIDE dataset [48] has two training sets: ITS with 13,990 indoor pairs and OTS with 313,950 outdoor pairs. The test set, SOTS, includes 500 indoor and 500 outdoor pairs. For real-world datasets, we use Dense-Haze [49] and NH-HAZE [50], each with 55 pairs. The last five pairs from each are used for testing, the rest for training.
Comparison settings. We compare DGFDNet with ten CNN-based methods: GridDehazeNet [24], MSBDN [41], FFA-Net [25], PMNet [42], MAXIM-2S [43], DEA-Net [28], FocalNet [38], OKNet [45], PGH2Net [47], and DCMPNet [46], along with three Transformer-based models: DeHamer [34], DehazeFormer [14], and Fourmer [44]. Performance is evaluated using PSNR and SSIM.
Training details. Each stage of the model consists of {2, 2, 4, 2, 2} DGFDBlocks with a base channel size of 32. We use the Adam optimizer ($\beta_{1}=0.9,\beta_{2}=0.999$) with a dataset-dependent initial learning rate, decaying to $1\times 10^{-6}$ via cosine annealing. FLOPs are calculated for $256\times 256$ patches. All experiments run on an NVIDIA 3090 GPU.
@@ IV-B1 Quantitative Comparisons
Table I quantitatively compares DGFDNet with SOTA methods on two synthetic and two real-world datasets. Bold and underlined values indicate the best and second-best results, respectively. DGFDNet outperforms all existing methods across all datasets.
Compared to traditional CNN-based methods like FocalNet [38] and OKNet [45], DGFDNet improves PSNR by 1.36 dB and 1.39 dB on SOTS-Indoor, and by 0.80 dB and 0.83 dB on SOTS-Outdoor, while using about half the parameters and less than half the FLOPs. Unlike FocalNet, which uses dual-domain selection with spatial attention, DGFDNet incorporates dark channel-guided attention for more accurate haze localization and a full-frequency modulation unit to better handle varying degradation levels.
Compared to the heavy DCMPNet [46], DGFDNet achieves comparable results on SOTS-Indoor with just 12.0% of the parameters and 19.7% of the FLOPs. On SOTS-Outdoor, it outperforms DCMPNet by 1.95 dB in PSNR and 0.002 in SSIM, showing better robustness in complex conditions. This improvement is driven by the PCGB module, which mitigates the limitations of the original dark channel prior. With similar resource budgets, DGFDNet consistently surpasses the lightweight PGH2Net [47], with 0.48 dB higher PSNR on SOTS-Indoor and 0.99 dB PSNR and 0.006 SSIM improvements on SOTS-Outdoor, balancing accuracy and efficiency.
On real-world datasets like Dense-Haze [49] and NH-HAZE [50], DGFDNet performs excellently. While its PSNR on NH-HAZE is slightly lower than Dehamer [34], it ranks second overall and achieves the highest SSIM on both datasets, demonstrating strong generalization and excellent global structure reconstruction in real-world hazy conditions.
@@ IV-B2 Qualitative Comparisons
Fig. 5 shows dehazing results on synthetic images from the SOTS dataset. Our method achieves the highest PSNR across all test images. In the first row (b, c), FFA-Net [25] and MAXIM-2S [43] produce blurred edges and artifacts in complex indoor scenes. MAXIM-2S [43] and Dehamer [34] also struggle with haze removal in deeper regions (2nd row, c, d). While FocalNet [38] and OKNet [45] perform better, our method still outperforms them. As shown in Fig. 5 (f), our approach removes haze more effectively and preserves sharp edges, achieving superior visual quality. In outdoor scenes, our method continues to outperform others, achieving better haze removal and clearer reconstruction.
Fig. 6 and Fig. 7 show real-world dehazing results on the Dense-Haze [49] and NH-HAZE [50] datasets. In Fig. 6, two dense haze scenes are compared. In the first, our method achieves slightly lower PSNR than Dehamer [34], but surpasses it by 0.183 in SSIM, offering clearer textures and reduced color distortion. In the second scene, with severe haze and significant detail loss, our method restores the overall structure and preserves more details than Dehamer [34].
Fig. 7 compares two non-homogeneous haze scenes. Apart from a slightly lower PSNR in the second scene compared to FocalNet [38], our method outperforms others in all metrics. In the first scene, it minimizes color distortion and avoids heavy artifacts from other methods. In the second scene, it removes haze most effectively, revealing clearer details, though it inadvertently eliminates some snow traces.
@@
9.79 dB 32.69 dB 34.27 dB 32.18 dB 33.67 dB 32.58 dB 35.20 dB PSNR
11.54 dB 34.49 dB 30.01 dB 34.38 dB 40.42 dB 40.29 dB 41.06 dB PSNR
22.04 dB 31.72 dB 29.04 dB 28.88 dB 35.11 dB 35.82 dB 39.72 dB PSNR
12.35 dB 26.62 dB 28.41 dB 26.86 dB 30.09 dB 31.83 dB 33.39 dB PSNR
Figure 5: Visual comparisons on synthetic hazy images from the SOTS dataset. Key regions highlighted by red boxes are enlarged in the lower-left corner for clearer comparison.
TABLE II: Ablation study for three core modules.
@@
IV-C Ablation Studies
We first evaluate the individual and combined contributions of HAFM, MGAM, and PCGB to verify their effectiveness and complementarity. Ablation studies are then conducted by modifying each module independently while keeping the rest of the architecture fixed. All models are trained on the ITS dataset and evaluated on SOTS-Indoor under the same conditions as the final model. Since PCGB targets the limitations of dark channel priors in outdoor scenes, we also validate its performance on SOTS-Outdoor using a model trained on the OTS dataset. Finally, attention map visualizations are provided to demonstrate the corrective effect of PCGB.
TABLE III: Ablation study for key components in HAFM.